OpenStudy (anonymous):

. Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions. x2 = 3x – 6

OpenStudy (anonymous):

\[ x^2 - 3 x +6 =0\\ a x^2 + b x + c=0\\ \Delta = b^2 - 4 a c \]

OpenStudy (anonymous):

\[\Delta = (-3)^2 - 4(1)(6)= 9 - 24= -15 \] What can you conclude?

OpenStudy (anonymous):

exactly one solution

OpenStudy (anonymous):

No. No real roots. Two complex roots \[ x=\frac{1}{2} \left(3-i \sqrt{15}\right) \\ x=\frac{1}{2} \left(3+i \sqrt{15}\right) \\ \]